# Distributed Optimal Coverage Control in Multi-agent Systems: Known and Unknown Environments

### v2

@@ IV-B Unknown Environment
where $r$ is a positive value applied during time $[0,\tau_{w}]$. The following theorem proves the convergence of the algorithm for unknown environments.
@@
Proof. Consider the following Lyapunov-like function:
@@
where $\tilde{\mathbf{A}}(t)\triangleq\{\tilde{\mathbf{a}}_{1}(t),\dots,\tilde{\mathbf{a}}_{n}(t)\}$. The time derivative of $V_{2}(\cdot)$ is as follows
$\dot{V}_{2}(\cdot)=\sum_{i=1}^{n}\Big{[}\frac{\partial\hat{H}^{\mathrm{T}}(t)}{\partial\mathbf{p}_{i}(t)}\ \!\dot{\mathbf{p}}_{i}(t)+\tilde{\mathbf{a}}_{i}^{\mathrm{T}}(t)\Gamma^{-1}\dot{\tilde{\mathbf{a}}}_{i}(t)\Big{]}.$ (45)
Using Lemma 3, and by substituting $\phi(\mathbf{q})$ from (35) and $\dot{\mathbf{p}}_{i}(t)$ from (23), one can write
$\dot{V}_{2}(\cdot)\!=$ $\!\sum_{i=1}^{n}\!\Big{[}\!-\!\int_{\mathbf{Q}}\!h_{\lambda}\!(t,R_{i})(\mathbf{q}-\mathbf{p}_{i}(t))^{\mathrm{T}}\hat{\phi}_{i}(\mathbf{q},t)\mathrm{d}\mathbf{q}$
$\epsilon\int_{Q}\!h_{\lambda}\!(t,R_{i})(\mathbf{q}-\mathbf{p}_{i}(t))\hat{\phi}_{i}(\mathbf{q},t)\mathrm{d}\mathbf{q}$
$+\epsilon\int_{\mathbf{Q}}\!h_{\lambda}\!(t,R_{i})\tilde{\mathbf{a}}_{i}^{\mathrm{T}}(t)\mathcal{K}(\mathbf{q})(\mathbf{q}-\mathbf{p}_{i}(t))^{\mathrm{T}}\mathrm{d}\mathbf{q}$
$\!\ \int_{\mathbf{Q}}\!h_{\lambda}\!(t,R_{i})(\mathbf{q}-\mathbf{p}_{i}(t))\hat{\phi}_{i}(\mathbf{q},t)\mathrm{d}\mathbf{q}$
@@
Finally, using (41) and by substituting $\dot{\hat{\mathbf{a}}}_{i}(t)$ from (37) in the above equation one can write
$\dot{V}_{2}(\cdot)\!=$ $\!\sum_{i=1}^{n}\!\Big{[}\!-\lVert\int_{\mathbf{Q}}h_{\lambda}\!(t,R_{i})(\mathbf{q}-\mathbf{p}_{i}(t))\hat{\phi}(\mathbf{q},t)\mathrm{d}\mathbf{q}\rVert^{2}$
$\qquad-\tilde{\mathbf{a}}_{i}^{\mathrm{T}}(t)\gamma\Lambda_{i}(t)\tilde{\mathbf{a}}_{i}(t)-\tilde{\mathbf{a}}^{\mathrm{T}}_{i}(t)\mathrm{\mathbf{I}}_{\mathrm{proj}_{i}}(t)\dot{\hat{\mathbf{a}}}_{\mathrm{pre}_{i}}(t)\Big{]}$
@@
For ease of referencing, denote the four terms in the right-hand side of (47) as $\varrho_{1}(t)$, $\varrho_{2}(t)$, $\varrho_{3}(t)$, and $\varrho_{4}(t)$, respectively. It follows that $\varrho_{4}(t)=-\zeta\sum_{j=1}^{m}\tilde{\mathbf{\Omega}}_{j}^{\mathrm{T}}(t)\mathbf{L}\hat{\mathbf{\Omega}}_{j}(t)$, where $\mathbf{\Omega}_{j}(t)\triangleq{a}^{j}\mathbf{1}_{n\times 1}$, $\hat{\mathbf{\Omega}}_{j}(t)\triangleq[\hat{{a}}_{1}^{j}(t),\dots,\hat{a}_{n}^{j}(t)]^{\mathrm{T}}$ and $\tilde{\mathbf{\Omega}}_{j}(t)\triangleq\hat{\mathbf{\Omega}}_{j}(t)-\mathbf{\Omega}_{j}(t)$. Based on the properties of the Laplacian matrix $\mathbf{L}$, $\mathbf{\Omega}_{j}^{\mathrm{T}}\mathbf{L}=a^{j}\mathbf{1}^{\mathrm{T}}_{n\times 1}\mathbf{L}=0,\forall j$. Therefore,
$-\zeta\sum_{j=1}^{m}\tilde{\mathbf{\Omega}}_{j}^{\mathrm{T}}(t)\mathbf{L}\hat{\mathbf{\Omega}}_{j}(t)=-\zeta\sum_{j=1}^{m}\hat{\mathbf{\Omega}}_{j}^{\mathrm{T}}(t)\mathbf{L}\hat{\mathbf{\Omega}}_{j}(t).$ (48)
It is clear that $\varrho_{1}(t)\leq 0$, and according to the definition of $w(t)$ and $\mathbf{L}$, $\varrho_{2}(t)$ and $\varrho_{4}(t)$ are also non-positive. Using (38), if $\hat{{a}}_{i}^{j}(t)>{a}_{\mathrm{min}}$ or $\hat{{a}}_{i}^{j}(t)={a}_{\mathrm{min}}$ and $\dot{\hat{{a}}}_{\mathrm{pre}_{i}}^{j}(t)\geq 0$, then $\mathrm{{I}}_{\mathrm{proj}_{i}}^{j}(t)=0$, and the corresponding term is vanished. On the other hand, if $\hat{{a}}_{i}^{j}(t)={a}_{\mathrm{min}}$ and $\dot{\hat{{a}}}_{\mathrm{pre}_{i}}^{j}(t)<0$, then $\mathrm{{I}}_{\mathrm{proj}_{i}}^{j}(t)=1$ and we observe that $\tilde{{a}}_{i}^{j}(t)=\hat{{a}}_{i}^{j}(t)-{a}^{j}\leq 0$, since ${a}^{j}\geq{a}_{\mathrm{min}}$. Therefore, $\varrho_{3}(t)\leq 0$, and, consequently, $\dot{V}_{2}(\cdot)\leq 0$. It can be shown that $\ddot{V}_{2}(\cdot)$ is bounded and $\dot{V}_{2}(\cdot)$ is therefore uniformly continuous. Since $V_{2}(\cdot)\geq 0$ (is lower bounded) and $\dot{V}_{2}(\cdot)\leq 0$, then $V_{2}(\cdot)$ has a finite limit. Therefore, using Barbalat’s lemma, $\dot{V}_{2}(\cdot)\to 0$, and as a result $\varrho_{i}(t)\to 0$, $i=1,\dots,4$. Hence it follows that:
@@
Additionally, based on Assumption 3, since $\dot{V_{2}}\to 0$ and $\lim_{t\to\infty}\Lambda_{i}(t)$ is positive definite, then $\tilde{\mathbf{a}}{i}\to 0$. Utilizing (35), it follows that $\lim_{t\to\infty}\tilde{\phi}_{i}(\mathbf{q},t)=0$. Thus, the estimation of the agent $i$ converges to the globally true approximation of the density function, $\phi(\mathbf{q})$. $\blacksquare$
@@ V Simulation Results
Value $0.1$ $2\times 10^{-3}$ $4$ $25\times 10^{-3}$ $0.2$
@@ V-B Unknown environments
Value $0.14$ $0.01$ $1$ $0.9$ $180$ $400\times\mathbf{1}_{25\times 1}$ $\mathbf{0}_{25\times 1}$ $0.2$
@@ V-C1 Complexity analysis
In the Voronoi coverage algorithm, the primary computation of the discrete control signal entails two main steps. First, there is a summation over all points in set $Q$, and then the computation of ${h}_{i}(\cdot)$ values. Calculating ${h}_{i}(\cdot)$ involves computing the distance from the point $q$ to all $n$ agents, which amounts to order $n$ operation, and comparing these distances between the current agent and the point to all other computed distances to determine the value of $h_{i}(\cdot)$, which involves $n-1$ comparisons. This results in a computation of complexity order $N(2n-1)$. Given that $N\gg 2n-1$, the complexity of this algorithm can be expressed as $\mathcal{O}(N)$.

### v3

@@ IV-B Unknown Environment
where $r$ is a positive value applied during time $0,\tau_{w}]$. The following theorem proves the convergence of the algorithm for unknown environments.
@@
Consider the following Lyapunov-like function:
@@
where $\tilde{\mathbf{A}}(t)\triangleq\{\tilde{\mathbf{a}}_{1}(t),\dots,\tilde{\mathbf{a}}_{n}(t)\}$. The time derivative of $V_{2}(\cdot)$ is as follows:
$\dot{V}_{2}(\cdot)=\frac{\partial\hat{H}}{\partial t}+\sum_{i=1}^{n}\Big{[}\frac{\partial\hat{H}^{\mathrm{T}}(t)}{\partial\mathbf{p}_{i}(t)}\ \!\dot{\mathbf{p}}_{i}(t)+\tilde{\mathbf{a}}_{i}^{\mathrm{T}}(t)\Gamma^{-1}\dot{\tilde{\mathbf{a}}}_{i}(t)\Big{]}.$ (49)
Using Lemma 3, one can write:
$\dot{V}_{2}(\cdot)\!=$ $\!\sum_{i=1}^{n}\!\Big{[}\frac{1}{2}\int_{Q}\frac{R_{i}\dot{\lambda}(t)}{\lambda^{2}(t)}h_{\lambda}\!(t,R_{i})\|\mathbf{q}-\mathbf{p}_{i}(t)\|^{2}\phi\left(\mathbf{q}\right)d\mathbf{q}$
$-\!\int_{\mathbf{Q}}\!\!\!h_{\lambda}\!(t,R_{i})(\mathbf{q}-\mathbf{p}_{i}(t))\!^{\mathrm{T}}(\hat{\phi}_{i}(\mathbf{q},t)-\tilde{\phi}_{i}(\mathbf{q},t))\dot{\mathbf{p}}_{i}(t)\mathrm{d}\mathbf{q}$
@@
Finally, using (45), and by substituting $\phi(\mathbf{q})$, $\dot{\mathbf{p}}_{i}(t)$ and $\dot{\hat{\mathbf{a}}}_{i}(t)$ from (39), (23) and (41), respectively, in the above equation one can write:
$\dot{V}_{2}(\cdot)\!$ $=\!\frac{1}{2}\sum_{i=1}^{n}\!\int_{Q}\frac{R_{i}\dot{\lambda}(t)}{\lambda^{2}(t)}h_{\lambda}(t,R_{i})\|\mathbf{q}-\mathbf{p}_{i}(t)\|^{2}\phi\!\left(\mathbf{q}\right)\!\mathrm{d}\mathbf{q}$
$+\sum_{i=1}^{n}\Big{[}-\lVert\int_{\mathbf{Q}}h_{\lambda}(t,R_{i})(\mathbf{q}-\mathbf{p}_{i}(t))\hat{\phi}_{i}(\mathbf{q},t)\mathrm{d}\mathbf{q}\rVert^{2}$
$-\tilde{\mathbf{a}}_{i}^{\mathrm{T}}(t)\gamma\Lambda_{i}(t)\tilde{\mathbf{a}}_{i}(t)-\tilde{\mathbf{a}}^{\mathrm{T}}_{i}(t)\mathrm{\mathbf{I}}_{\mathrm{proj}_{i}}\!(t)\dot{\hat{\mathbf{a}}}_{\mathrm{pre}_{i}}\!(t)\Big{]}$
@@
For ease of referencing, denote the five terms in the right-hand side of (51) as $\varrho_{1}(t)$, $\varrho_{2}(t)$, $\varrho_{3}(t)$, $\varrho_{4}(t)$, and $\varrho_{5}(t)$, respectively. It follows that $\varrho_{5}(t)=-\zeta\sum_{j=1}^{m}\tilde{\mathbf{\Omega}}_{j}^{\mathrm{T}}(t)\mathbf{L}\hat{\mathbf{\Omega}}_{j}(t)$, where $\mathbf{\Omega}_{j}(t)\triangleq{a}^{j}\mathbf{1}_{n\times 1}$, $\hat{\mathbf{\Omega}}_{j}(t)\triangleq[\hat{{a}}_{1}^{j}(t),\dots,\hat{a}_{n}^{j}(t)]^{\mathrm{T}}$ and $\tilde{\mathbf{\Omega}}_{j}(t)\triangleq\hat{\mathbf{\Omega}}_{j}(t)-\mathbf{\Omega}_{j}(t)$. Based on the properties of the Laplacian matrix $\mathbf{L}$, $\mathbf{\Omega}_{j}^{\mathrm{T}}\mathbf{L}=a^{j}\mathbf{1}^{\mathrm{T}}_{n\times 1}\mathbf{L}=0,\forall j$. Therefore,
$\zeta\sum_{j=1}^{m}\tilde{\mathbf{\Omega}}_{j}^{\mathrm{T}}(t)\mathbf{L}\hat{\mathbf{\Omega}}_{j}(t)=-\zeta\sum_{j=1}^{m}\hat{\mathbf{\Omega}}_{j}^{\mathrm{T}}(t)\mathbf{L}\hat{\mathbf{\Omega}}_{j}(t).$ (52)
It is clear that $\varrho_{1}(t)<0$, $\varrho_{2}(t)\leq 0$, and according to the definition of $w(t)$ and $\mathbf{L}$, $\varrho_{3}(t)$ and $\varrho_{5}(t)$ are also non-positive. Using (42), if $\hat{{a}}_{i}^{j}(t)>{a}_{\mathrm{min}}$ or $\hat{{a}}_{i}^{j}(t)={a}_{\mathrm{min}}$ and $\dot{\hat{{a}}}_{\mathrm{pre}_{i}}^{j}(t)\geq 0$, then $\mathrm{{I}}_{\mathrm{proj}_{i}}^{j}(t)=0$, and the corresponding term is vanished. On the other hand, if $\hat{{a}}_{i}^{j}(t)={a}_{\mathrm{min}}$ and $\dot{\hat{{a}}}_{\mathrm{pre}_{i}}^{j}(t)<0$, then $\mathrm{{I}}_{\mathrm{proj}_{i}}^{j}(t)=1$ and we observe that $\tilde{{a}}_{i}^{j}(t)=\hat{{a}}_{i}^{j}(t)-{a}^{j}\leq 0$, since ${a}^{j}\geq{a}_{\mathrm{min}}$. Therefore, $\varrho_{4}(t)\leq 0$, and, consequently, $\dot{V}_{2}(\cdot)<0$. It can be shown that $\ddot{V}_{2}(\cdot)$ is bounded and $\dot{V}_{2}(\cdot)$ is therefore uniformly continuous. Since $V_{2}(\cdot)>0$ (is lower bounded) and $\dot{V}_{2}(\cdot)<0$, then $V_{2}(\cdot)$ has a finite limit. Therefore, using Barbalat’s lemma, $\dot{V}_{2}(\cdot)\to 0$, and as a result $\varrho_{i}(t)\to 0$, $i=1,\dots,5$. Hence, it follows that:
@@
Additionally, based on Assumption 3, since $\dot{V_{2}}\to 0$ and $\lim_{t\to\infty}\Lambda_{i}(t)$ is positive definite, then $\tilde{\mathbf{a}}_{i}\to 0$. Utilizing (39), it follows that $\lim_{t\to\infty}\tilde{\phi}_{i}(\mathbf{q},t)=0$. Thus, the estimation of the agent $i$ converges to the true approximation of the density function, $\phi(\mathbf{q})$. $\blacksquare$
@@ V Simulation Results
Value $0.1$ $2\times 10^{-3}$ 44 $25\times 10^{+3}$ $0.2$
@@ V-B Unknown environments
Value $0.14$ $0.01$ 11 $0.9$ $180$ $400\times\mathbf{1}_{25\times 1}$ $\mathbf{0}_{25\times 1}$ $0.2$
@@ V-C1 Complexity analysis
In the Voronoi coverage algorithm, the primary computation of the discrete control signal entails two main steps. First, there is a summation over all points in set $Q$, and then the computation of ${h}_{i}(\cdot)$ values. Calculating ${h}_{i}(\cdot)$ involves computing the distance from the point $q$ to all $n$ agents, which amounts to order $n$ operation, and comparing these distances between the current agent and the point to all other computed distances to determine the value of $h_{i}(\cdot)$, which involves n1n-1 comparisons. This results in a computation of complexity order N(2n1)N(2n-1). Given that $N\gg 2n-1$, the complexity of this algorithm can be expressed as $\mathcal{O}(N)$.